OpenStudy (anonymous):

Can someone help me? thankss Find the equation of the the tangent line to the curve: y=1 + xe^2x

OpenStudy (thomas5267):

At which point?

OpenStudy (anonymous):

x=0

OpenStudy (anonymous):

Take the derivative of the function, evaluate at x=0; that will be the slope of your line (m). Use the point/slope form of the equation of the line by using that value of m and (0,f(0)).

OpenStudy (anonymous):

Alright thank you, but another question. How would you solve for something like this:

OpenStudy (anonymous):

No such (real) value of x exists. Exponential functions are strictly positive.

OpenStudy (anonymous):

Okay but in general, how woul you solve an equation with e?

OpenStudy (campbell_st):

use the product rule to find the derivative u = x du/dx = 1 v = e^2x dv/dx = 2e^2x then \[dy/dx = 2xe^{2x} + e^{2x} = e^{2x}( 2x + 1)\] now substitute x = 0 to find the gradient of the tangent \[m = 2^{2\times0}(2\times0 + 1) = 1\] the tangent has a gradient of 1 at x = 0 the point on the curve when x = 0 is y = 1 so you have a gradient of m = 1 and point (0,1) then the tangent is y = 1 = 1(x - 1) y = x + 1 is the equation of the tangent at x = 0

OpenStudy (anonymous):

thank you all for your answers :D